OpenStudy (anonymous):

PLZ HELP...... if an object is dropped from a height of 144 feet, the function h(t)=-16t^2+144 gives the hight of the object after t seconds. when will the object hit the ground. A.9s B.6s C.1.5s D.3s

OpenStudy (anonymous):

"hit the ground" means h(t)=0. Solve \(144-16t^2=0\).

OpenStudy (anonymous):

sorry the 16 is -

OpenStudy (anonymous):

So... \(16t^2=144\) What might t be? (You can ignore negative values.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

Have you ever solved a quadratic equation before?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Then reason your way through it. You have perfect squares on each side of the equals sign, so take the square root of each side.

OpenStudy (anonymous):

square of 16 is 4 and square of 144 is 12

OpenStudy (anonymous):

Yep, so what's left?

OpenStudy (anonymous):

4 and 12

OpenStudy (anonymous):

would you divid the two numbers

OpenStudy (anonymous):

Yes, you need to get t by itself on one side of the equals sign.

OpenStudy (anonymous):

so then it would be three

OpenStudy (anonymous):

Yep. Three what? What is the question asking for?

OpenStudy (anonymous):

You can also verify the solution in the original function: \(h(t)=-16t^2+144 \space \rightarrow h(3)=-16(3)^2+144=0.\)

OpenStudy (anonymous):

oh thank you very much for your help